# Miniaturized optical system for a chip based cold atom inertial sensor

### v1

@@ I Introduction
To use cold Rubidium atom technologies for atomic clocks in space Aveline et al. (2020); Laurent et al. (2020), for on board gravimetry Bidel et al. (2018, 2020), or microgravity experiments Geiger et al. (2011); Müntinga et al. (2013); Rudolph et al. (2015); Becker et al. (2018) a lot of effort has been conducted to develop and ruggedized the laser system needed to cool and manipulate them. As an example, some mobile gravimeters use laser diodes in extended cavities at 780 nm with semiconductor amplifiers Schmidt et al. (2011a, b); Cheinet (2006); Merlet et al. (2014); Zhang et al. (2018) with a free space optical system for routing and switching all the laser beams needed in the device. This approach has been ruggedized to operate in space Laurent et al. (2006); Lévèque et al. (2015); Elliott et al. (2018), a drop tower and a sounding rocket Schiemangk et al. (2015); Schkolnik et al. (2016); Dinkelaker et al. (2017); Pahl et al. (2019); Strangfeld et al. (2021) where a volume of 43 liters has been achieved Pahl et al. (2019). Another approach is to use frequency doubling of telecom lasers at 1.56 $\mu$m Lienhart et al. (2007); Carraz et al. (2009); Theron et al. (2015, 2017); Diboune et al. (2017); Dingjan et al. (2006); Nyman et al. (2006); Stern et al. (2009); Ménoret et al. (2011). This allows higher power at 780 nm Sané et al. (2012) than with direct emission at 780 nm using diode lasers and semiconductor amplifiers and the whole system can be fully fiber-based Lévèque et al. (2014); Legg and Farries (2017); Diboune et al. (2017).
@@ II.2 The miniaturized optical system
The frequency used for the cooling corresponds to the transition $\left|5^{2}S_{1/2},F=2\right>\rightarrow\left|5^{2}P_{3/2},F=3\right>$ of the hyperfine structure of Rubidium 87 for D2 line as shown in Figure 2. The cooling of an atom works on a closed cycling transition. However, this one is not closed because the frequency difference between $\left|5^{2}P_{3/2},F=3\right>$ and $\left|5^{2}P_{3/2},F=2\right>$ is only about 200 MHz, thus some atoms can go to $\left|5^{2}P_{3/2},F=2\right>$ leading to a possible decay in $\left|5^{2}S_{1/2},F=1\right>$. After a large number of absorption-emission cycles, all atoms will fall into that state and there will be no longer atoms in the cooling cycle. Therefore, we need a second laser, the pumping laser that is resonant with the $\left|5^{2}S_{1/2},F=1\right>\rightarrow\left|5^{2}P_{3/2},F=2\right>$ transition, to recycle atoms fallen into the $\left|5^{2}S_{1/2},F=1\right>$ state back into the cooling cycle. The two inputs of the bench are for the two cooling and pumping lasers, together cooling the atoms. The cooling laser is tuned to the cooling transition and the pumping laser is tuned to the pumping transition (see Figure 2).
Inside the optical bench, Polarization Beam Splitters (PBS) allows for separation and recombination of the beams. Each output needed for the realization of the 2DMOT and 3DMOT (MOT 3DX1, MOT 3DX2, MOT 3DH, MOT 2D/PushBeam) has the two frequencies. Acousto-Optical Modulators (AOM) in the bench allow for adjusting the laser frequencies. We can shift the frequency of the cooling beam during the optical molasses (using AOM4 which shifts the cooling frequency before sending it to the saturated absorption spectroscopy), the optical pumping (using AOM6) and atom detection (using AOM1). For those shifts, we use AOMs in a double-pass configuration Donley et al. (2005). The pumping laser beam is locked on the transition between $\left|5^{2}S_{1/2},F=1\right>$ and the crossover between $\left|5^{2}P_{3/2},F=1\right>$ and $\left|5^{2}P_{3/2},F=2\right>$ using saturated absorption spectroscopy, which is not the pumping frequency, thus the AOM3 in the bench shifts the laser to the pumping frequency. Adjustable half wave plates let us adjust the optical powers on each output. Mechanical shutters extinguish beams during the interferometry sequence and assure that no light arrives to the atoms.
@@ III.2 Pumping laser frequency lock
Figure 6: (Color online) Saturated absorption signal is in blue (this signal amplitude is 30 times amplified for better reading) for the $|5^{2}S_{1/2},F=1>$ state of 87Rb atoms. Error signal is in orange. The zero corresponds to 384.234 604 760 THz.
@@ III.3 Cooling laser frequency lock
The vacuum cell of our sensors is made up of two chambers as described in Dupont-Nivet (2016); Wirtschafter (2022), the first one for the 2DMOT, the second one for the 3DMOT. Both are connected through a differential pumping aperture of 750 $\mu$m diameter. With two pairs of counter propagating laser beams, the PushBeam and a magnetic field gradient created by magnets, we realize a 2DMOT to cool atoms in the xy plane (see Figure 8.b) in the first chamber. The magnets are designed to create a magnetic field gradient in the plane xy and invariant along the z axis with a zero in the center of the cell where the atoms will accumulate. We obtain a thin beam of atoms which velocity direction is oriented along the z axis. The PushBeam pushes the atoms from the first chamber to the second chamber along the z axis through the differential pumping aperture. Then, with three pairs of counter propagating laser beams and a magnetic field gradient along the x, y and z axis created by two coils in anti-Helmoltz configuration, we realize a 3DMOT in the second chamber Gajda and Mostowski (1994); Devlin and Tarbutt (2016). The realization of a 2DMOT before a 3DMOT allows charging rapidly a larger number of atoms in the 3DMOT, while keeping the vacuum below $10^{-10}$ mbar in the second chamberFarkas et al. (2010). In the 3DMOT and 2DMOT, the counter propagating laser beams are created with mirrors on which the beam reflect and overlap with the original beam.
Figure 9: (Color online) (a) 2DMOT and (b) 3DMOT obtained using the beams prepared by the miniaturized optical bench. By fluorescence, we see the atoms in the 3DMOT. Atom number in the 3DMOT is about $2.5\cdot 10^{8}$.

### v2

# Miniaturized optical system for a chip-based cold-atom inertial sensor

@@ I Introduction
To use cold Rubidium atom technologies for atomic clocks in space Aveline et al. (2020); Laurent et al. (2020), for on board gravimetry Bidel et al. (2018, 2020), or microgravity experiments Geiger et al. (2011); Müntinga et al. (2013); Rudolph et al. (2015); Becker et al. (2018) a lot of effort has been conducted to develop and ruggedized the laser system needed to cool and manipulate atoms. As an example, some mobile gravimeters use laser diodes in extended cavities at 780 nm with semiconductor amplifiers Schmidt et al. (2011a, b); Cheinet (2006); Merlet et al. (2014); Zhang et al. (2018) and with a free space optical system for routing and switching all the laser beams needed in the device. This approach has been ruggedized to operate in space Laurent et al. (2006); Lévèque et al. (2015); Elliott et al. (2018), a drop tower and a sounding rocket Schiemangk et al. (2015); Schkolnik et al. (2016); Dinkelaker et al. (2017); Pahl et al. (2019); Strangfeld et al. (2021) where a volume of 43 liters has been achieved Pahl et al. (2019). Another approach is to use frequency doubling of telecom lasers at 1.56 $\mu$m Lienhart et al. (2007); Carraz et al. (2009); Theron et al. (2015, 2017); Diboune et al. (2017); Dingjan et al. (2006); Nyman et al. (2006); Stern et al. (2009); Ménoret et al. (2011). This allows higher power at 780 nm Sané et al. (2012) than with direct emission at 780 nm using diode lasers and semiconductor amplifiers and the whole system can be fully fiber-based Lévèque et al. (2014); Legg and Farries (2017); Diboune et al. (2017). A non-exhaustive review of these systems is presented in table 1.
@@ II.2 The miniaturized optical system
The frequency used for the cooling corresponds to the transition $\left|5^{2}S_{1/2},F=2\right>\rightarrow\left|5^{2}P_{3/2},F=3\right>$ of the hyperfine structure of Rubidium 87 for D2 line as shown in Figure 2. The cooling of an atom works on a closed cycling transition. However, this one is not closed because the frequency difference between $\left|5^{2}P_{3/2},F=3\right>$ and $\left|5^{2}P_{3/2},F=2\right>$ is only about 266 MHz, thus some atoms can go to $\left|5^{2}P_{3/2},F=2\right>$ leading to a possible decay in $\left|5^{2}S_{1/2},F=1\right>$. After a large number of absorption-emission cycles, all atoms will fall into that state and there will be no longer atoms in the cooling cycle. Therefore, we need a second laser, the pumping laser that is resonant with the $\left|5^{2}S_{1/2},F=1\right>\rightarrow\left|5^{2}P_{3/2},F=2\right>$ transition, to recycle atoms fallen into the $\left|5^{2}S_{1/2},F=1\right>$ state back into the cooling cycle. The two inputs of the bench are for the two cooling and pumping lasers, together cooling the atoms. The cooling laser is tuned to the cooling transition and the pumping laser is tuned to the pumping transition (see Figure 2).
Inside the optical bench, Polarization Beam Splitters (PBS) allows for separation and recombination of the beams. Each output needed for the realization of the 2DMOT and 3DMOT (MOT 3DX1, MOT 3DX2, MOT 3DH, MOT 2D/PushBeam) has the cooling and pumping frequencies. Acousto-Optical Modulators (AOM) in the bench allow for adjusting the laser frequencies. We can shift the frequency of the cooling beam during the optical molasses (using AOM4 which shifts the cooling frequency before sending it to the saturated absorption spectroscopy), the optical pumping (using AOM6) and atom detection (using AOM1). For those shifts, we use AOMs in a double-pass configuration Donley et al. (2005). The pumping laser beam is locked on the transition between $\left|5^{2}S_{1/2},F=1\right>$ and the crossover between $\left|5^{2}P_{3/2},F=1\right>$ and $\left|5^{2}P_{3/2},F=2\right>$ using saturated absorption spectroscopy, which is not the pumping frequency, thus the AOM3 in the bench shifts the laser to the pumping frequency. Adjustable half wave plates let us adjust the optical powers on each output. Mechanical shutters extinguish beams during the interferometry sequence and assure that no light arrives to the atoms.
@@ III.2 Pumping laser frequency lock
Figure 6: (Color online) Saturated absorption signal is in blue (this signal is 30 times amplified for better reading) for the $|5^{2}S_{1/2},F=1>$ state of 87Rb atoms. Error signal is in orange. The zero corresponds to 384.234 604 760 THz.
@@ III.3 Cooling laser frequency lock
Figure 9: (Color online) (a) 2DMOT and (b) 3DMOT obtained using the beams prepared by the miniaturized optical bench. By fluorescence, we see the atoms in the 2DMOT and 3DMOT. Atom number in the 3DMOT is about $2.5\cdot 10^{8}$. Axis are the same than figure 8.
The vacuum cell of our sensors is made up of two chambers as described in Dupont-Nivet (2016); Wirtschafter (2022), the first one for the 2DMOT, the second one for the 3DMOT. Both are connected through a differential pumping aperture of 750 $\mu$m diameter. With two pairs of counter propagating laser beams, the PushBeam and a magnetic field gradient created by magnets, we realize a 2DMOT to cool atoms in the xy plane in the first chamber (see Figure 8.b). The magnets are designed to create a magnetic field gradient in the plane xy and invariant along the z axis with a zero in the center of the cell where the atoms will accumulate. We obtain a thin beam of atoms which velocity direction is oriented along the z axis. The PushBeam pushes the atoms from the first chamber to the second chamber along the z axis through the differential pumping aperture. Then, with three pairs of counter propagating laser beams and a magnetic field gradient along the x, y and z axis created by two coils in anti-Helmoltz configuration, we realize a 3DMOT in the second chamber Gajda and Mostowski (1994); Devlin and Tarbutt (2016) (see Figure 8.a). The realization of a 2DMOT before a 3DMOT allows charging rapidly a larger number of atoms in the 3DMOT, while keeping the vacuum below $10^{-10}$ mbar in the second chamberFarkas et al. (2010). In the 3DMOT and 2DMOT, the counter propagating laser beams are created with mirrors on which the beam reflect and overlap with the original beam.